# IMPACT: Behavioral Intention-aware Multimodal Trajectory Prediction with Adaptive Context Trimming

### v3

@@ 4.2 Leaderboard Performance
Marginal Prediction Performance.
Table 3 shows that our method achieves state-of-the-art (SOTA) performance on the Waymo 2024 motion prediction benchmark, outperforming all existing LiDAR-free approaches in the primary metric, Soft mAP, as well as in minADE and Overlap Rate. In the remaining metrics (mAP, minFDE, and Miss Rate), our method ranks second. Here, ’Ensemble’ denotes the model-ensemble technique described in [shi2023MTR] for performance boosting, while ’Single’ indicates evaluation with a single model.
@@
As presented in Table 4, even without any model-ensemble techniques [shi2023MTR], our single model achieves the best performance on the Waymo joint prediction leaderboard, surpassing the previous SOTA method [liu2024reasoningmultiagentbehavioraltopology] by 10.2% in both Soft mAP and mAP. Furthermore, our model attains the second-lowest Overlap Rate and the third-lowest Miss Rate. These substantial improvements underscore the effectiveness of the IMPACT framework.
4.3 Behavioral Intention Prediction Performance
@@ 4.4 Vectorized Occupancy Prediction Performance
Table 2 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results are in Appendix.
RMP-YOLO(Ensemble)[sun2024rmpyolorobustmotionpredictor] 0.4737 0.4531 0.5564 1.1188 0.1084 ${0.1259}^{*}$
ModeSeq(Ensemble)[zhou2024modeseq] 0.4737 0.4665 0.5680 1.1766 0.1204 0.1275
BeTop[liu2024reasoningmultiagentbehavioraltopology] 0.4698 0.4587 0.5716 1.1668 0.1183 0.1272
MGTR[gan2024mgtrmultigranulartransformermotion] 0.4599 0.4505 0.5918 1.2135 0.1298 0.1275
EDA[lin2024eda] 0.4596 0.4487 0.5718 1.1702 0.1169 0.1266
MTR++[shi2024mtr++] 0.4410 0.4329 0.5906 1.1939 0.1298 0.1281
MTR[shi2023MTR] 0.4403 0.4249 0.5964 1.2039 0.1312 0.1274
HDGT[JiaHDGT] 0.3709 0.3577 0.5933 1.2055 0.1511 0.1557
DenseTNT[gu2021densetnt] - 0.3281 1.0387 1.5514 0.1573 0.1779
SceneTransformer[ngiam2022scenetransformer] - 0.2788 0.6117 1.2116 0.1564 0.1473
Ours (Ensemble) 0.4801 ${0.4598}^{*}$ 0.5563 1.1295 0.1087 0.1258
Ours (Single) 0.4721 0.4609 ${0.5641}^{*}$ ${1.1540}^{*}$ ${0.1143}^{*}$ 0.1255
BeTop-ens [liu2024reasoningmultiagentbehavioraltopology] 0.2573 0.2511 0.9779 2.2805 0.4376 $0.1688^{*}$
BeTop [liu2024reasoningmultiagentbehavioraltopology] $0.2466^{*}$ $0.2412^{*}$ 0.9744 2.2744 0.4355 0.1696
MTR++ [shi2024mtr++] 0.2368 0.2326 $\underline{0.8975}$ $\underline{1.9509}$ 0.4143 0.1665
MTR [shi2023MTR] 0.2078 0.2037 $0.9181^{*}$ 2.0633 0.4411 0.1717
MotionDiffuser [jiang2023motiondiffuser] 0.2047 0.1952 0.8642 1.9482 $\underline{0.4300}$ 0.2004
GameFormer [huang2023gameformergametheoreticmodelinglearning] 0.1982 0.1923 0.9721 2.2146 0.4933 0.2022
BIFF [zhu2023biffbilevelfuturefusion] 0.1383 0.1229 0.9388 $2.0443^{*}$ 0.4620 0.2172
DenseTNT [gu2021densetnt] - 0.1647 1.1417 2.4904 0.5350 0.2309
M2I [sun2022m2ifactoredmarginaltrajectory] - 0.1239 1.3506 2.8325 0.5538 0.2757
SceneTransformer [ngiam2022scenetransformer] - 0.1192 0.9774 2.1892 0.4942 0.2067
Ours (Single) 0.2718 0.2659 0.9738 2.2734 $0.4316^{*}$ 0.1684
4.5 Ablation Study
Different behavior modeling and map interaction mechanisms.
Table 5 presents an ablation study evaluating various agent-pruning and map-pruning strategies on the validation set. The first row (All + Dynamic) queries all agents for the decoder while employing dynamic map selection as in MTR [shi2023MTR], serving as a baseline. Introducing agent selection via Braid Theory [liu2024reasoningmultiagentbehavioraltopology] (Row 2) enhances Soft mAP to 0.4602. Replacing Braid Theory with our two-class filtering—where Yielding, Overtaking, and Nearby behaviors are treated as interactive while Ignore is considered non-interactive (Row 3)—further refines minADE to 0.5693 while maintaining a comparable Miss Rate. Notably, adopting our four-class YOIN approach (Row 4) elevates Soft mAP to 0.4728, demonstrating its ability to better capture nuanced interactions. Lastly, incorporating vectorized occupancy into YOIN (Row 5) further improves performance (Soft mAP = 0.4758), highlighting the benefit of explicit map filtering in isolating relevant polylines and refining trajectory predictions. These results validate the effectiveness of our integrated behavior modeling and map selection framework.
Agent Map Metrics
All w/Braid Theory[liu2024reasoningmultiagentbehavioraltopology] w/Inter. w/YOIN w/Dynamic[shi2023MTR] w/Vect. Occ. Soft mAP mAP minADE Miss Rate Inference Time (ms/scenario) Params (M)
0.4582 0.4414 0.5718 0.1203 14.67 69.469
0.4602 0.4507 0.5745 0.1187 23.92 48.938
0.4663 0.4549 0.5693 0.1187 23.92 48.938
0.4728 0.4613 0.5680 0.1163 23.92 48.941
0.4758 0.4646 0.5652 0.1173 18.27 48.944
@@ 4.5 Ablation Study
On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 5 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 5) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[liu2024reasoningmultiagentbehavioraltopology] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations. Notably, the approach retains or improves accuracy metrics such as Soft mAP, mAP, and minADE, underscoring that targeted context selection effectively balances computational efficiency with predictive performance.
Table 6: Performance gains from integrating proposed modules with existing approaches (Argoverse 1 & 2). bFDE6 is the main metric.

### v4

@@ 4.3 Ablation Study
As presented in Table 2, even without any model-ensemble techniques [27], our single model achieves the best performance on the Waymo joint prediction leaderboard, surpassing the previous SOTA method [21] by 10.2% in both Soft mAP and mAP. Furthermore, our model attains the second-lowest Overlap Rate and the third-lowest Miss Rate. These substantial improvements underscore the effectiveness of the IMPACT framework.
@@ 4.5 Vectorized Occupancy Prediction Performance
Table 5 presents precision, recall, and F1-score for both classes, along with their data ratios in the validation set. The occupied class constitutes only 3.96% of samples, highlighting a strong imbalance. Despite this, the model achieves a high overall accuracy of 0.989. For the occupied class, precision (0.933) and recall (0.775) yield an F1-score of 0.847, indicating rare false positives—crucial for pruning irrelevant map polylines in trajectory prediction. The dominant unoccupied class achieves near-perfect precision (0.991), recall (0.998), and F1-score (0.994). These results demonstrate that our vectorized occupancy prediction effectively identifies critical map segments while minimizing false classification. More qualitative results (like Figure 4) are in Appendix section 7.
@@
As shown in Figure 6, selecting 24 agents and 192 polylines achieves the optimal balance between interaction diversity and trajectory accuracy. On average, the model input contains 43.85 agents and 749.52 map polylines, meaning this selection reduces scene complexity by 45.2% and 74%, respectively, while preserving essential contextual cues. Beyond this threshold, redundant elements introduce noise, degrading generalization and increasing prediction uncertainty. This highlights the importance of selective context pruning to enhance trajectory forecasting performance.
@@
We compare various agent-pruning and map-pruning strategies in Table 3 using the same encoder architecture. The first row (All + w/Dynamic) corresponds to a baseline that queries all agents and dynamically collect polylines, resulting in a model with 69.469 M parameters and a reported inference time of 14.67 ms under a simplified setting. By contrast, our final approach (the last row in Table 3) combines YOIN-based agent pruning with vectorized occupancy-based pruning. This design lowers the parameter count to 48.944 M while reducing the average inference time of BeTOP[21] from 23.92 to 18.27 ms per scenario. The key to these efficiency gains lies in pruning irrelevant scene information—both agents and polylines—based on their semantic contributions, which in turn eliminates redundant computations.
Table 6: Performance gains from integrating proposed modules with existing approaches(Argoverse1&2). bFDE6 is main metric.
@@ 5 Conclusion
Supplementary Material
1 Behavioral Intention Categories Split.
The selection of four intention labels: nearby, ignore, overtaking, and yielding—is motivated by balancing model expressiveness, computational efficiency, and actionable interpretability for downstream decision-making.
  1. 1.
    Limitations of Oversimplified Labels (Interactive/Ignore). While a binary classification reduces complexity, it fails to capture nuanced interaction types, agents labeled as ”interactive” may exhibit vastly different kinematic patterns (e.g., merging vs. lane-keeping), which cannot be resolved without finer-grained labels.
  2. 2.
    Drawbacks of Overly Complex Labels. Complex taxonomies struggle to generalize across diverse interaction patterns. Also, high-dimensional label spaces dilute feature representations, reducing inter-class separability. Lastly, rare categories lack sufficient training samples, causing models to memorize noise rather than learn generalizable patterns.
  3. 3.
    The rationality of the four categories. According to algorithm2, these four labels have corresponding interaction information, which covers most interaction modes in urban driving. 4D embeddings also achieve optimal trade-offs between class separation and parameter efficiency.
2 Training Loss Detail
The overall training objective comprises four components: $\mathcal{L}_{\mathrm{Int}},\mathcal{L}_{\mathrm{Occ}},\mathcal{L}_{\mathrm{Traj}}$ and $\mathcal{L}_{\mathrm{Score}}$. Since $\mathcal{L}_{\mathrm{Int}},\mathcal{L}_{\mathrm{Occ}},\mathcal{L}_{\mathrm{Traj}}$ are multimodal to avoid mode-collapse, we apply winner-takes-all strategy. The weighting factors are set as $\lambda_{\mathrm{1}}=100$, $\lambda_{\mathrm{2}}=100$, $\lambda_{\mathrm{3}}=1$, and $\lambda_{\mathrm{4}}=1$.
2.1 Behavioral Intention Prediction.
The behavioral intention module predicts $\hat{H}^{(k)}\in\mathbb{R}^{N_{a}\times 4}$ for each modality $k$ and each agent, and the four output classes represent different intentions (yielding, overtaking, ignore, and nearby). The ground-truth label $H^{*}\in\mathbb{R}^{N_{a}\times 4}$ is a one-hot encoding of the true class for each agent. Once the winning modality $k^{*}$ is determined, we use $\hat{H}^{(k^{*})}$ to compute the Focal Loss [20] in a multi-class setting:
$\mathcal{L}_{\mathrm{int}}=-\sum_{n=1}^{N_{a}}\sum_{c=1}^{4}\alpha_{c}\Bigl(1-\hat{H}_{n,c}^{(k^{*})}\Bigr)^{\gamma_{c}}\,H_{n,c}^{*}\,\log\bigl(\hat{H}_{n,c}^{(k^{*})}\bigr),$ (16)
where $\alpha_{c}=[0.1,0.45,0.45,0.45]$ and $\gamma_{c}=[2,1,1,1]$ are class-dependent balancing and focusing parameters, respectively. Intuitively, $\alpha_{c}$ compensates for class imbalance (i.e., classes with fewer training examples receive higher weight), while $\gamma_{c}$ “focuses” more strongly on hard, misclassified examples.
2.2 Vectorized Occupancy Prediction.
In the vectorized occupancy module, the prediction $\hat{O}\in\mathbb{R}^{N_{l}\times 1}$ represents the probability that each of the $N_{l}$ map polylines is occupied by the target agent in the future. The ground-truth label is $O^{*}\in\{0,1\}^{N_{l}\times 1}$. We again use only the winning modality’s output $\hat{O}^{(k^{*})}$ and compute a binary Focal Loss:
$\mathcal{L}_{\mathrm{occ}}=$ $-\alpha\sum_{l=1}^{N_{l}}\Bigl[O_{l}^{*}\Bigl(1-\hat{O}_{l}^{(k^{*})}\Bigr)^{\gamma}\log\Bigl(\hat{O}_{l}^{(k^{*})}\Bigr)$ (17)
$+\bigl(1-O_{l}^{*}\bigr)\Bigl(\hat{O}_{l}^{(k^{*})}\Bigr)^{\gamma}\log\bigl(1-\hat{O}_{l}^{(k^{*})}\bigr)\Bigr].$
where $\alpha=0.25$ and $\gamma=2$. This choice of Focal Loss mitigates the class imbalance between “occupied” and “not occupied” labels and focuses the training on hard-to-predict polylines.
2.3 GMM-based Trajectory Loss.
Given the predicted future Gaussian components $\hat{\mathrm{Y}}\in\mathbb{R}^{K\times T_{f}\times 5}$, each modality $k$ predicts parameters $\bigl(\mu_{x}^{(k)},\mu_{y}^{(k)},\sigma_{x}^{(k)},\sigma_{y}^{(k)},\rho^{(k)}\bigr)$ for a bivariate Gaussian distribution at each future timestep $t$. Let the ground-truth future trajectories be $\mathrm{Y}^{*}=\{(x_{t}^{*},y_{t}^{*})\}_{t=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times 2}$. We compute the Negative Log-Likelihood (NLL) loss only for the selected (winning) modality $k^{*}$:
$\mathcal{L}_{\mathrm{traj}}=-\sum_{t=1}^{T_{f}}\log\Bigl(\mathcal{N}\bigl(\mathbf{x}_{t}^{*};\,\boldsymbol{\mu}_{t}^{(k^{*})},\boldsymbol{\Sigma}_{t}^{(k^{*})}\bigr)\Bigr),$ (18)
where $\boldsymbol{\mu}_{t}^{(k^{*})}=[\mu_{x,t}^{(k^{*})},\mu_{y,t}^{(k^{*})}]$ and covariance matrix $\boldsymbol{\Sigma}_{t}^{(k^{*})}$ is defined as:
$\boldsymbol{\Sigma}_{t}^{(k^{*})}=\begin{bmatrix}\sigma_{x,t}^{(k^{*})2}&\rho_{t}^{(k^{*})}\sigma_{x,t}^{(k^{*})}\sigma_{y,t}^{(k^{*})}\\ \rho_{t}^{(k^{*})}\sigma_{x,t}^{(k^{*})}\sigma_{y,t}^{(k^{*})}&\sigma_{y,t}^{(k^{*})2}\end{bmatrix}$. (19)
2.4 Mode Classification Loss.
Additionally, each modality $k$ predicts a confidence score $\hat{S_{k}}\in[0,1]$. We wish to train these scores so that the winning modality receives higher scores while others receive lower scores. Let $\delta^{*}\in\{0,1\}^{K}$ be a label vector that has $\delta_{k^{*}}^{*}=1$ for the winning modality and $\delta_{k}^{*}=0$ for all others. We use the standard binary cross-entropy:
[…]
Figure 7: Visualization of multimodal prediction results for the target agent, along with the top-1 mode behavioral intention predictions of surrounding agents. Ground-truth trajectories are also visualized to facilitate clear validation of the predicted behaviors.
Figure 8: Visualization of predicted multi mode intention labels. The first column renders the ground truth intention labels of the agents. Remaining columns render the predicted results of K=6 different modes. We can see that our predicted intentions and trajectories are coupled and cover most possible modes.
Figure 9: Visualization of ground truth behaviour intention labels comparison between BeTop(Covered, Ignored) and IMPACT(Nearby, Yielding, Overtaking, Ignored). The first row of the figures renders the result of BeTop.
Figure 10: Visualization of predicted multi modes vectorized occupancy. The first column renders the ground truth vectorized occupancy along the agents’ future trajectories. Remaining columns render the predicted results of K=6 different modes.
Figure 11: Visualization of multimodal predicted vectorized occupancy for the target agents in different scenarios. The darker the polylines, the greater the probability of occupation.
3 Model Ensemble Details.
Given $N$ well-trained models, each model produces 64 distinct future trajectories, resulting in a total of $64N$ multimodal trajectories for each target agent. Each trajectory is assigned a confidence score by its originating model.
First, we apply a softmax normalization to the confidence scores of all $64N$ trajectories. Next, we perform non-maximum suppression (NMS) based on the endpoints of these trajectories to select the top six. The NMS distance threshold $\sigma$ scales according to the length $L$ of the highest-confidence trajectory among the $64N$ predictions, as follows:
$\sigma=\min\left(3.5,\max\left(2.5,\frac{L-10}{50-10}\times 1.5+2.5\right)\right).$ (21)
The performance of IMPACT(Ensemble) is generated by an ensemble of four different models:
  1. 1.
    IMPACT model.
  2. 2.
    IMPACT without Behavioral Intention Prediction.
  3. 3.
    IMPACT without IMPACT Occupancy Prediction.
  4. 4.
    IMPACT with QCNet refinement module.
4 Network Details
[…]
5 Autolabel Algorithms for Behavioral Intention and Vectorized Occupancy
The autolabeling algorithms consist of two parts, Vectorized GT Occupancy Label Generation 1 and GT behavioral Intention Label Generation 2.
Algorithm 1 Vectorized GT Occupancy Label Generation
0:   Map polyline points $L\in\mathbb{R}^{N_{l}\times N_{p}\times 2}$;    Target agent’s future trajectory $y\in\mathbb{R}^{T_{f}\times 2}$;    Distance threshold $\alpha$;
0:   Occupancy labels $O\in\{0,1\}^{N_{l}\times 1}$;
1:  Broadcast $L$ to $N_{l},\,N_{p},\,1,\,2\,]$ and $y$ to [ $1,\,1,\,T_{f},\,2\,]$.
2:  Compute the distance tensor $d\in\mathbb{R}^{N_{l}\times N_{p}\times T_{f}}$.
3:  $O=(\min(d,\text{axis}=1)\leq\alpha\;)\text{.any}(\text{axis}=-1)$.
4:  return $O\in\mathbb{R}^{N_{l}\times 1}$
Algorithm 2 GT behavioral Intention Label Generation
0:Agents’ future trajectories $Y^{*}\in\mathbb{R}^{N_{a}\times T_{f}\times 2}$; Target agent’s future trajectory $y\in\mathbb{R}^{T_{f}\times 2}$;
0:    behavioral Intention labels $H\in\mathbb{R}^{N_{a}\times 4}$;
1:  Radius Screening:
2:  for $i=1\dots N_{a}$ do
3:   Compute $d_{i}(t)=\|Y^{*}_{i,t}-y_{t}\|_{2},\;\forall\,t=1\dots T_{f}.$
4:   if $\min_{t}\,d_{i}(t)>\tau_{\mathrm{ignore}}$ then
5:    Label $i$ as ignored and exclude from further checks.
6:   else
7:    Add $i$ to the candidate set $\mathcal{C}$.
8:   end if
9:  end for
10:  Interaction Detection:
11:  for $i\in\mathcal{C}$ do
12:   Define geometry $S_{i}(t)$ for agent $i$, and $S_{\mathrm{target}}(t)$ for the target agent.
13:   If $\Bigl(\forall\,t:\;S_{i}(t)\cap S_{\mathrm{target}}(t)=\varnothing\Bigr)\,\wedge\,\bigl(\min_{t}d_{i}(t)\leq\tau_{\mathrm{ignore}}\bigr)$:    label agent $i$ as nearby.
14:   Else if $\exists\,t:\;S_{i}(t)\cap S_{\mathrm{target}}(t)\neq\varnothing$:    compare the time steps of minimal distances to assign either overtaking or yielding.
15:  end for
16:  Label Assignment:
17:  Convert each final label {ignored, nearby, overtaking, yielding} into a one-hot vector $H_{i}\in\mathbb{R}^{4}$.
18:  return $H$
6 Per-Category Marginal and Joint Prediction
As shown in 8 and 9, we present a comprehensive per-category evaluation of our framework’s performance on both single-agent (marginal) and multi-agent (joint) motion prediction tasks using the Waymo Open Motion Dataset[39].
7 Additional Qualitative Results
Additional prediction results. Figure 12 provides qualitative prediction results under both joint and marginal settings.
Visualization results of multimodal behavior intention prediction. Figure 7 and 8 provide detailed forecast visualization of multimodal trajectories and behavior intention labels.
Behavior intention labels comparison between BeTop and IMPACT. Figure 9 provides ground truth behavior intention labels comparison between BeTop and IMPACT.
Visualization results of multimodal vectorized occupancy. Figure 11 and 10 provide detailed forecast visualization of multimodal trajectories and vectorized occupancy prediction.
8 Deployment Detail
We re-trained our method on a large-scale private dataset comprising 8.7 million clips for training and 1 million clips for validation. Each clip includes one second of historical data and eight seconds of future data. Our method is integrated into an end-to-end planning framework, in which it consumes upstream perception outputs—such as object detection, tracking, and map reconstruction—and generates trajectory predictions for the downstream planning module. We have successfully deployed our system on real vehicles, and a demonstration video is provided in the supplementary materials.
9 Notations
As shown in 10, we provide a lookup table for notations in the paper.
Table 7: Detailed performance of marginal predictions on WOMD Motion Leaderboard.
Our Model Category Soft mAP $\uparrow$ mAP $\uparrow$ minADE $\downarrow$ minFDE $\downarrow$ Miss rate $\downarrow$ Overlap Rate $\downarrow$
IMPACT Vehicle 0.5040 0.4897 0.6696 1.3672 0.1122 0.0430
Pedestrian 0.4963 0.4866 0.3394 0.7009 0.0644 0.2651
Cyclist 0.4159 0.4063 0.6832 1.3940 0.1662 0.0684
Avg 0.4721 0.4609 0.5641 1.1540 0.1143 0.1255
IMPACT (Ensemble) Vehicle 0.5035 0.4766 0.6668 1.3584 0.1111 0.0398
Pedestrian 0.4996 0.4887 0.3386 0.7011 0.0637 0.2664
Cyclist 0.4371 0.4142 0.6635 1.3290 0.1513 0.0713
Avg 0.4801 0.4598 0.5563 1.1295 0.1087 0.1258
IMPACT (e2e) Vehicle 0.4859 0.4657 0.6157 1.2309 0.1194 0.0464
Pedestrian 0.4599 0.4389 0.3114 0.6360 0.0716 0.2668
Cyclist 0.3842 0.3712 0.6419 1.2822 0.1911 0.0774
Avg 0.4434 0.4253 0.5230 1.0497 0.1274 0.1302
Table 8: Marginal predictions on WOMD Motion Leaderboard
Category Method minADE minFDE Miss Rate mAP Soft mAP
Vehicle MTR[27] 0.7642 1.5257 0.1514 0.4494 0.4590
ModeSeq[47] 0.678 1.4046 0.1129 0.5095 0.5181
BeTopNet[21] 0.6814 1.3888 0.1172 0.4860 0.4995
IMPACT 0.6668 1.3584 0.1111 0.4766 0.5035
Pedestrian MTR[27] 0.3486 0.7270 0.0753 0.4331 0.4409
ModeSeq[47] 0.3407 0.7117 0.0776 0.4709 0.4781
BeTopNet[21] 0.3451 0.7142 0.0668 0.4777 0.4875
IMPACT 0.3386 0.7011 0.0637 0.4887 0.4996
Cyclist MTR[27] 0.7022 1.4093 0.1786 0.3561 0.3650
ModeSeq[47] 0.6852 1.4136 0.1706 0.4190 0.4248
BeTopNet[21] 0.6905 1.3975 0.1688 0.4060 0.4163
IMPACT 0.6635 1.3290 0.1513 0.4142 0.4371
Table 9: Joint predictions on WOMD Interaction Leaderboard
Category Method minADE minFDE Miss Rate mAP Soft mAP
Vehicle GameFormer[14] 1.0499 2.4044 0.4321 0.2469 0.2564
BeTopNet[21] 1.0216 2.3970 0.3738 0.3374 0.3308
IMPACT 1.0267 2.4162 0.3713 0.3380 0.3450
Pedestrian GameFormer[14] 0.7978 1.8195 0.4713 0.1962 0.2014
BeTopNet[21] 0.7862 2.3970 0.3738 0.3374 0.3308
IMPACT 0.7834 1.8198 0.4005 0.2849 0.2917
Cyclist GameFormer[14] 1.0686 2.4199 0.5765 0.1367 0.1338
BeTopNet[21] 1.1155 2.5850 0.5253 0.1717 0.1756
IMPACT 1.1113 2.5841 0.5231 0.1748 0.1788
Figure 12: Qualitative Results of IMPACT for Joint and Marginal Prediction on the WOMD Validation Dataset.
Table 10: List of important notations in this work
Symbol Meaning Dimension
$N_{a}$ Number of agents in the scene Scalar
$N_{l}$ Number of polylines in the map Scalar
$N_{p}$ Number of points in each polyline Scalar
$T_{p}$ Number of past timesteps (observation horizon) Scalar
$T_{f}$ Number of future timesteps (prediction horizon) Scalar
$K$ Number of future modes (prediction branches) Scalar
$M$ Number of query content features/modes in the decoder Scalar
$D$ Feature dimension in the latent space Scalar
$\mathcal{A}=\{a_{1},\dots,a_{N_{a}}\}$ Set of all agent trajectories (history) $\mathbb{R}^{N_{a}\times T_{p}\times F_{1}}$
$\mathcal{L}=\{l_{1},\dots,l_{N_{l}}\}$ Set of all map polylines $\mathbb{R}^{N_{l}\times N_{p}\times F_{2}}$
$\mathcal{R}=\{r_{1},\dots,r_{N_{l}}\}$ Historical relative movement features $\mathbb{R}^{N_{l}\times T_{p}\times F_{3}}$
$y_{i}^{k}$ $k$-th predicted future trajectory for agent $i$ $\mathbb{R}^{T_{f}\times 2}$
$s_{i}^{k}$ Confidence score for $k$-th future trajectory Scalar
$\mathcal{H}^{k}=\{h_{1}^{k},\dots,h_{N_{a}}^{k}\}$ behavioral intentions under mode $k$ $\mathbb{R}^{N_{a}\times 4}$
$\mathcal{O}^{k}=\{o_{1}^{k},\dots,o_{N_{l}}^{k}\}$ Occupancy probabilities under mode $k$ $\mathbb{R}^{N_{l}\times 1}$
$\mathcal{A}_{k}^{\mathrm{sel}}$ Top-$m$ selected agents under mode $k$ $\mathbb{R}^{m\times D}$
$\mathcal{L}_{k}^{\mathrm{sel}}$ Top-$n$ selected polylines under mode $k$ $\mathbb{R}^{n\times D}$
$\phi(\cdot),\varphi(\cdot)$ Utility functions deriving scalar scores from vectors
$Q$ Query content features in the decoder $\mathbb{R}^{M\times D}$
$\hat{\mathrm{O}}$ Predicted vectorized occupancy $\mathbb{R}^{K\times N_{l}\times 1}$
$\mathrm{O}^{*}$ Ground-truth vectorized occupancy $\mathbb{R}^{N_{l}\times 1}$
$\hat{\mathrm{H}}$ Predicted behavioral intentions $\mathbb{R}^{K\times N_{a}\times 4}$
$\mathrm{H}^{*}$ Ground-truth behavioral intentions $\mathbb{R}^{N_{a}\times 4}$
$\hat{\mathrm{Y}}$ Predicted Future Gaussian components $\mathbb{R}^{K\times T_{f}\times 5}$
$\mathrm{Y}^{*}$ Ground-truth Future Trajectories $\mathbb{R}^{T_{f}\times 2}$
$\hat{\mathrm{S}}$ Predicted scores $\mathbb{R}^{K}$